OpenStudy (anonymous):

given that the x intercept of a line is (-21/2,0) and the y intercept is (0,5) what is the equation of the line in slope intercept form?

OpenStudy (anonymous):

(y-5)/x=(-5/-21/2)=10/21

OpenStudy (anonymous):

no that does not seem to be correct the answer should end up as y =

OpenStudy (anonymous):

Slope-intercept form is y=mx+b. You already have the y-intercept (b) so all you need to do is find the slope (m). To find the slope, you take the change in y and divide it by the change in x. So, the change in y is 5-0=0. The change in x is 0-(-21/2) = 21/2. Be careful, when finding the slope, you must use the parts of the order pair in the same order. Otherwise, your sign will be wrong. So the slope is \[5/(21/2)=10/21\] Plugging these back into the form, we get \[y=(10/21)x + 5\]

OpenStudy (anonymous):

thats the wrong answer i need the right one !!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

How do you know it's wrong? I got the same slope as brackett's post, and the y-intercept was given.

OpenStudy (anonymous):

what is the y intercept

OpenStudy (anonymous):

the point where it crosses the y-axis (0,y). Your problem gave it as (0,5). It's also the value of b in the slope-intercept form.

OpenStudy (anonymous):

so what is the answer to my problem

OpenStudy (anonymous):

The last line of my original post.

OpenStudy (anonymous):

its not the right answer acoording to the computer

OpenStudy (anonymous):

Well, all I can say is that the computer is wrong. Unless someone else can tell me where I messed up. I've done thousands of problems just like this. Are you typing the answer into a program that tells you it's wrong? If so, it might be that the way it expects the answer (namely the fraction) to be is different than they way you're typing (or pasting) it.

OpenStudy (anonymous):

ok

OpenStudy (shadowfiend):

Right you were, mtbender74. Well explained.

OpenStudy (anonymous):

thank you.